# Deep Learning of Force Manifolds from the Simulated Physics of Robotic Paper Folding

### 05 Feb 2024 v3

@@ VI-A Dual Manipulator Setup
For our experiments, we use two Rethink Robotics’ Sawyer manipulators as shown in Fig. 7. One arm has an elongated gripper designed for folding, while the other arm has a spring-compliant roller for creasing and an Intel Realsense D435 camera for visual feedback. The elongated gripper has rubber attached to the insides of the fingers for tight gripping.
@@ VI-B Perception System
VI-C Visual-Feedback Control
Although we minimize $\lambda$ with our proposed framework, sliding could still occur due to a substrate’s low friction surface and/or jittering of the robot’s end-effector. Notice that the optimal trajectory $\tau^{*}$ generated as described in Sec. V-D assumes that the origin $\mathbf{o}$ of our coordinate system, shown in Fig. 4(a), is fixed. We can define the origin as $\mathbf{o}=\mathbf{q}_{0}-l\hat{\mathbf{x}}$, where $l$ is the total length of the paper. Any amount of sliding indicates that $\mathbf{q}_{0}$ is moving along the $x$-axis and, therefore, the origin $\mathbf{o}$ also moves an identical amount. When this occurs, our position within the manifold during traversal deviates from the optimal trajectory. Furthermore, without adaptive replanning, the amount of sliding $\Delta x$ will directly result in $\Delta x$ amount of error when creasing. To circumvent this, we introduce a visual-feedback approach that mitigates the effects of sliding.
@@ VI-C Visual-Feedback Control
We then replan a partial trajectory $\tau^{*}_{i}$ from the updated start state $(x^{c}_{i},z_{i})$ to the next state $(x_{i+1},z_{i+1})$ in the sequence and carry out this updated trajectory. This is repeated until reaching the goal state. By properly accounting for sliding, we ensure that the traversal through the NFM is as accurate as possible. Note that this scheme allows us to obtain corrected partial trajectories in near real time once $N$ becomes sufficiently large, as each partial trajectory’s goal state approaches its start state, allowing for uniform cost search to conclude rapidly. We direct the reader to our supplementary videos (Footnote 1), which showcase the speed of the feedback loop.
@@ VII-A Measuring the Material Property of Paper
where $l_{h}$ is the vertical distance from the free end to the fixed end and $l$ is the total length of the paper. We can obtain the mapping $\mathcal{L}(\epsilon)$ using numerical simulations, which is shown in Fig. 9(b). With this mapping known, simple algebra can be performed to obtain $L_{gb}$. First, we measure the ratio $\epsilon=l_{h}/l$ for a particular paper to obtain its corresponding normalized total length $\bar{l}$. Then, the value of $L_{gb}$ can be calculated simply by $L_{gb}=l/\bar{l}$. Once we obtain $L_{gb}$, we can now use the non-dimensionlized mapping (11) to find the optimal path for manipulating the paper.
@@ VII-B Baseline Algorithms
$\tau_{B}$ =$=\{(R\cos(i\textrm{d}\theta),R\sin(i\textrm{d}\theta),i\textrm{d}\theta)\ \forall\ i\in[0,M]\},$
@@
Additionally, we conducted comparisons against the state-of-the-art mechanics-based folding algorithm presented by Petrik et al. [15, 17], which we refer to as the “SOTA baseline”, that uses a beam model to compute folding trajectories for fabric minimizing sliding. However, this baseline considers only the current status of the deformed material when computing subsequent optimal grasp and, consequently, is unable to handle the challenging task of paper folding. Examples of the computed trajectories are shown in Fig. 10.
Algorithm A4 $\mathcal{C}=0.25$ m A4 $\mathcal{C}=$ Half US $\mathcal{C}=0.20$ m US $\mathcal{C}=$ Half CA $\mathcal{C}=0.20$ m CA $\mathcal{C}=$ Half OR $\mathcal{C}=0.30$ m
Petrik et al.[15, 17] $59.46$ $51.15$ $68.14$ $47.30$ $80.07$ $77.28$ $43.20$
Our Method $3.28$ $4.13$ $1.80$ $2.28$ $1.27$ $4.19$ $11.73$
@@ VII-C Experimental Setup
For the rectangular papers (1–3), we performed two sets of experiments. The first involved folding the papers to an arbitrary crease location ($C=0.25$ m for A4 and $C=0.20$ m for US Letter and cardboard), while the second involves folding the papers in half. For the square origami paper, we chose an arbitrary crease location of $C=0.30$ m. This resulted in a total of 7 folding scenarios. For each of the scenarios, we conducted experiments using 4 different algorithms—the intuitive baseline, the SOTA baseline, our open-loop approach, and our closed-loop approach. We completed 10 trials for each of these algorithms, resulting in a total of 280 experiments.
@@ VII-E Parameters
For closed-loop control, we chose to split all trajectories into $N=5$ intervals regardless of trajectory length. Furthermore, we used a slick (i.e., low friction) table to showcase the robustness of our method. Note that smaller friction coefficients result in a significantly harder manipulation problem due to the lower threshold for sliding. This is made evident by the excessive sliding of the baseline algorithms shown later. Using an empirical method, we conducted measurements to determine the static coefficient of friction between the papers and substrate, yielding an approximate value of $\mu_{s}=0.12$. For comparison, the static coefficient of friction for lubricated steel on steel is typically $\mu_{s}=0.15$.
@@ VII-F Results and Analysis
In Table I we report the offline trajectory computation times for all experiments using a single Intel i9-9900KF CPU, demonstrating on average a 15×15$\times$15 × speed improvement over the SOTA baseline. In Fig. 11, all experimental results are reported as box plots where we show the achieved fold lengths and spin errors. From the achieved fold lengths, we see significant improvement over the two baselines for all folding scenarios. As expected, the SOTA baseline demonstrates better performance compared to the intuitive method with the exception of cardboard paper where the former fails to fold it at all. Due to the large gap in performance, broken axes are used to properly display the variance of the recorded data. Note that not only do our algorithms achieve significantly better performance on average, the variance of our approach is also much lower as shown by the decreased $y$-axis resolution after the axis break. We attribute the high variances of the baseline algorithms to the increased influence of friction, which can often cause chaotic, unpredictable results. In other words, truly deterministic folding can only be achieved when sliding is nonexistent.
@@
The sliding problem is only exacerbated by increasing the stiffness of the material ($L_{gb}=0.132$). Fig. 13(a) shows the intuitive baseline algorithm failing to fold the cardboard paper in half by a margin almost as long as the paper itself, while Fig. 13(b) shows how the SOTA baseline method experiences complete failure due to a high energy snapping caused by excessive deformation. By comparison, our open-loop algorithm is capable of folding the cardboard with significantly better results albeit with some sliding, as shown in Fig. 13(c). As the material stiffness increases, the benefits of visual feedback are more clearly seen as we are able to achieve near perfect folding for cardboard, as shown in Fig. 13(d). All of our findings for rectangular folding also match the results of our diagonal folding experiment shown in Fig. 12(b1)–(b3), where the closed-loop approach once again achieves minimal sliding when compared to the baselines. Overall, the matching findings across all our experiments demonstrate the robustness of our formulation against material and geometric factors.
@@
In terms of spin error, we found that softer materials had the greatest error. As the surface of the table is not perfectly flat, any amount of sliding will directly result in uneven spin as shown in Fig. 12(a). As the material stiffness increases, the spin errors became more uniform across the methods as the influence of friction is not enough to deform the paper. Nevertheless, we can see that our open and closed-loop algorithms resulted in less sliding than the baseline on average.
VIII Conclusion
@@
We have introduced a novel robot control strategy capable of robustly folding sheets of paper of varying materials and geometries with only a single manipulator. Our framework incorporates a combination of techniques spanning several disciplines, including physical simulation, machine learning, scaling analysis, and path planning. The effectiveness of the framework was demonstrated through extensive real world experiments against both natural and state-of-the-art paper folding strategies. Furthermore, an efficient, near real-time visual feedback algorithm was implemented that further minimizes folding error. Our closed-loop sensorimotor control algorithm successfully accomplished challenging scenarios such as folding stiff cardboard with repeatable accuracy.
@@
In future work, we hope to tackle the difficult problem of creating arbitrary creases along sheets of paper with non-symmetric centerlines. Such non-symmetric papers can no longer be represented as a reduced-order 2D elastic rod model, hence requiring a more sophisticated formulation. Additionally, folding along regions of paper with preexisting creases and folds will be a crucial step to achieving elaborate folding tasks, such as robotic origami. Moving forward, we anticipate exploring solutions to such problems that take advantage of generalized problem formulations with data-driven control schemes such as reinforcement learning.
Acknowledgments
We acknowledge financial support from the National Science Foundation under Grant numbers IIS-1925360, CAREER-2047663, and OAC-2209782.

### 20 Feb 2024 v4

@@ VI Robotic System
Figure 7: Example of our perception system with a top-down view of the folding procedure. (a) Shows the intuitive baseline results while (b) shows our open-loop algorithm for $L_{gb}=0.048$ and $C=0.25$ m. As in Fig. 2, the solid green line indicates the desired end effector position while the dashed blue line indicates the crease location. For this case, we observe that the intuitive baseline suffers from considerable sliding while our open-loop algorithm has near-perfect performance.
@@ VI-A Dual Manipulator Setup
For our experiments, we use two Rethink Robotics’ Sawyer manipulators, as shown in Fig. 6. One arm has an elongated gripper designed for folding, while the other arm has a spring-compliant roller for creasing and an Intel Realsense D435 camera for visual feedback. The elongated gripper has rubber attached to the insides of the fingers for tight gripping.
@@ VI-B Perception System
VI-C Model-Predictive Control via Visual Feedback
Although we minimize $\lambda$ with our proposed framework, sliding could still occur due to a substrate’s low friction surface and/or jittering of the robot’s end-effector. Notice that the optimal trajectory $\tau^{*}$ generated as described in Sec. V-D assumes that the origin $\mathbf{o}$ of our coordinate system, shown in Fig. 4(a), is fixed. We can define the origin as $\mathbf{o}=\mathbf{q}_{0}-l\hat{\mathbf{x}}$, where $l$ is the total length of the paper. Any amount of sliding indicates that $\mathbf{q}_{0}$ is moving along the $x$-axis and, therefore, the origin $\mathbf{o}$ also moves an identical amount. When this occurs, our position within the manifold during traversal deviates from the optimal trajectory. Furthermore, without adaptive replanning, the amount of sliding $\Delta x$ will directly result in $\Delta x$ amount of error when creasing. To circumvent this, we introduce a model-predictive control approach that mitigates the effects of sliding through trajectory corrections via visual feedback.
@@ VI-C Model-Predictive Control via Visual Feedback
We then replan a partial trajectory $\tau^{*}_{i}$ from the updated start state $(x^{c}_{i},z_{i})$ to the next state $(x_{i+1},z_{i+1})$ in the sequence and carry out this updated trajectory. This is repeated until reaching the goal state. By properly accounting for sliding, we ensure that the traversal through the NFM is as accurate as possible. Note that this scheme allows us to obtain corrected partial trajectories in near real time once $N$ becomes sufficiently large, as each partial trajectory’s goal state approaches its start state, allowing for uniform cost search to conclude rapidly. We refer the reader to our supplementary videos (Footnote 1), which showcase the speed of the feedback loop.
@@
Input: $(x_{s},z_{s}),(x_{g},z_{g}),L_{gb},\delta,N,\mathcal{F}_{\textrm{NN}}$
crease paper with roller
Algorithm 2 Closed-loop Control Pseudocode
@@ VII-A Measuring the Material Property of Paper
where $l_{h}$ is the vertical distance from the free end to the fixed end and $l$ is the total length of the paper. We can obtain the mapping $\mathcal{L}(\epsilon)$ using numerical simulations (Fig. 9(b)). With this mapping known, simple algebra can be performed to obtain $L_{gb}$. First, we measure the ratio $\epsilon=l_{h}/l$ for a particular paper to obtain its corresponding normalized total length $\bar{l}$. Then, the value of $L_{gb}$ can be calculated simply by $L_{gb}=l/\bar{l}$. Once we obtain $L_{gb}$, we can now use the non-dimensionlized mapping (11) to find the optimal path for manipulating the paper.
@@ VII-B Baseline Algorithms
$\tau_{B}$ =$=\{(R\cos(i\textrm{d}\theta),R\sin(i\textrm{d}\theta),i\textrm{d}\theta),\ \forall i\in[0,M]\},$
@@
Additionally, we conducted comparisons against the state-of-the-art mechanics-based folding algorithm presented by Petrik et al. [3, 5], to which we refer as the “SOTA baseline”, which uses a beam model to compute folding trajectories for fabric minimizing sliding. However, this baseline considers only the current status of the deformed material when computing subsequent optimal grasp and, consequently, is unable to handle the challenging task of paper folding. Examples of the computed trajectories are shown in Fig. 10.
@@ VII-C Experimental Setup
For the rectangular papers (1–3), we performed two sets of experiments. The first involved folding the papers to an arbitrary crease location ($C=0.25$ m for A4 and $C=0.20$ m for US Letter and cardboard), while the second involves folding the papers in half. For the square origami paper, we chose an arbitrary crease location of $C=0.30$ m. This resulted in a total of 7 folding scenarios. For each of the scenarios, we conducted experiments using 4 different algorithms—the intuitive baseline, the SOTA baseline, our open-loop approach, and our closed-loop approach. We completed 10 trials for each of these algorithms, resulting in 280 experiments.
We also validated our model’s non-dependence on the paper’s width $w$ (Sec. IV-A) through additional experiments involving narrow strip folding. We created strips of width $w=2.5$ cm for both A4 and cardboard and performed 10 trials for each algorithm, resulting in 80 additional experiments for a total of 360 experiments in our extensive case study.
Algorithm A4 $C=0.25$ m A4 $C=$ Half US $C=0.20$ m US $C=$ Half CA $C=0.20$ m CA $C=$ Half OR $C=0.30$ m
Petrik et al.[3, 5] $59.46$ $51.15$ $68.14$ $47.30$ $80.07$ $77.28$ $43.20$
Our Method $3.28$ $4.13$ $1.80$ $2.28$ $1.27$ $4.19$ $11.73$
@@
TABLE II: Evaluation of the Influence of Paper Width
Paper $\left(C/2\right)$ $w$ [cm] Fold length $\textrm{mean}\pm\sigma$ [cm] Spin error $\theta_{\textrm{err}}$ $\textrm{mean}\pm\sigma$ [$\deg$]
INT-base SOTA-base Open-loop Closed-loop INT-base SOTA-base Open-loop Closed-loop
A4 (14.85) $2.5$ $10.89\pm 0.69$ $14.66\pm 0.46$ $14.88\pm 0.07$ $14.83\pm 0.04$ $0.36\pm 1.65$ $0.02\pm 0.65$ $0.48\pm 0.47$ $-0.23\pm 0.48$
$21$ $12.31\pm 0.63$ $13.51\pm 0.13$ $14.74\pm 0.05$ $14.64\pm 0.07$ $6.56\pm 1.29$ $1.68\pm 0.97$ $0.27\pm 0.44$ $0.35\pm 0.38$
CB (14.0) $2.5$ $7.00\pm 0.45$ N/A $13.31\pm 0.14$ $13.81\pm 0.08$ $0.07\pm 0.75$ N/A $-0.70\pm 0.31$ $-0.35\pm 0.82$
$21$ $5.88\pm 0.38$ N/A $12.8\pm 0.27$ $13.75\pm 0.19$ $1.25\pm 1.43$ N/A $1.38\pm 1.06$ $1.45\pm 1.48$
@@ VII-E Parameters
For closed-loop control, we chose to split all trajectories into $N=5$ intervals regardless of trajectory length. Furthermore, we used a slick (i.e., low friction) table to demonstrate the robustness of our method. Note that smaller friction coefficients result in a significantly harder manipulation problem due to the lower threshold for sliding. This is made evident by the excessive sliding of the baseline algorithms shown later. Using an empirical method, we conducted measurements to determine the static coefficient of friction between the papers and substrate, yielding an approximate value of $\mu_{s}=0.12$.
@@ VII-F Results and Analysis
In Table I we report the offline trajectory computation times for all experiments using a single Intel i9-9900KF CPU, demonstrating on average a 15×15$\times$15 × speed improvement over the SOTA baseline. In Fig. 11, all experimental results for non-narrow strips are reported as box plots where we show the achieved fold lengths and spin errors. From the achieved fold lengths, we see significant improvement over the two baselines for all folding scenarios. As expected, the SOTA baseline demonstrates better performance compared to the intuitive method with the exception of cardboard paper where the former fails to fold it at all. Due to the large gap in performance, broken axes are used to properly display the variance of the recorded data. Note that not only do our algorithms achieve significantly better performance on average, the variance of our approach is also much lower as shown by the decreased $y$-axis resolution after the axis break. We attribute the high variances of the baseline algorithms to the increased influence of friction, which can often cause chaotic, unpredictable results. In other words, truly deterministic folding can only be achieved when sliding is nonexistent.
@@
The sliding problem is only exacerbated by increasing the stiffness of the material ($L_{gb}=0.132$). Fig. 13(a) shows the intuitive baseline algorithm failing to fold the cardboard paper in half by a margin almost as long as the paper itself, while Fig. 13(b) shows how the SOTA baseline method experiences complete failure due to a high energy snapping caused by excessive deformation. By comparison, our open-loop algorithm is capable of folding the cardboard with significantly better results albeit with some sliding (Fig. 13(c)). As the material stiffness increases, the benefits of visual feedback are more clearly seen as we are able to achieve near perfect folding for cardboard (Fig. 13(d)). All of our findings for rectangular folding also match the results of our diagonal folding experiment shown in Fig. 12(b1)–(b3), where the closed-loop approach once again achieves minimal sliding when compared to the baselines. Overall, the matching findings across all our experiments demonstrate the robustness of our formulation against material and geometric factors.
@@
In terms of spin error, we found that softer materials had the greatest error. As the surface of the table is not perfectly flat, any amount of sliding will directly result in uneven spin, as shown in Fig. 12(a). As the material stiffness increases, the spin errors became more uniform across the methods as the influence of friction is not enough to deform the paper. Nevertheless, we can see that our open and closed-loop algorithms resulted in less sliding than the baseline on average.
VII-G Effects of Paper Width on Folding Performance
Previously, we mathematically deduced that the paper’s width $w$ should have no influence on our folding scheme (Sec. IV-A). We now validate this claim through experiments. Table II reports comparisons of fold length and spin error between narrow strips and wide sheets of paper error for half folding A4 paper and cardboard. As expected, both our open-loop and closed-loop methods have near identical performance regardless of paper width. Aside from a slight exception for open-loop narrow cardboard folding that actually benefits from a $5$ mm improvement, differences are almost imperceptible to the human eye. Interestingly, both baseline methods experience significant changes in fold length (>±$>\pm 1$ cm) when width is changed. The prevention of such nondeterministic behavior is another benefit of our method.
VIII Additional Discussion
VIII-A Performing Multiple Folds on the Same Paper
Our optimal folding strategy can also be used to fold a single piece of paper multiple times so long as our assumptions regarding material homogeneity and symmetrical centerline hold. Given our method’s exceptional accuracy, a precise equilateral origami cube can be created using solely open-loop control, and excellent results are shown in Fig. 14. Future work pertaining to performing arbitrary folds is discussed in Sec. IX.
VIII-B Importance of Single-Manipulator Folding
Figure 14: Folding a simple origami cube with open-loop control. The left image shows the starting strip of paper (A4), the middle images show three folding steps to create each corner, and the right image shows the resulting cube. Note that regrasps were performed manually.
Figure 15: Isometric views for folding $C=\textrm{Half}$ with the stiffest paper ($L_{gb}=0.132$) using an auxiliary manipulator: (a) shows the intuitive baseline while (b) shows the SOTA baseline. Note the similarity in results to Fig. 13 despite the use of an auxiliary manipulator to prevent sliding.
To emphasize the significance of single-manipulator folding, we next present results for half-folding cardboard using both baselines with the incorporation of an auxiliary manipulator used to prevent the opposing edge from sliding (Fig. 15). The auxiliary manipulator “holds down” the edge during the folding process and then moves out of the way at the last possible point in the folding trajectory in order to avoid collision. Despite the additional hardware, we can see that the results for both baselines are near identical to the original experiments shown in Fig. 13.
For the intuitive baseline, significant sliding occurs as soon as the auxiliary manipulator lifts away due to the paper’s high energy state, ultimately yielding an end result that is nowhere near the desired half-fold. This sliding is less noticeable for the SOTA baseline, but the sheet still suffers from buckling. With this demonstration, we conclude that for stiff materials, naive folding strategies can fail drastically despite the use of auxiliary manipulators to explicitly prevent sliding. In fact, the use of auxiliary manipulators is a rather expensive approach, both in terms of requiring additional hardware as well as introducing more overall motions into the pipeline, which our folding strategy completely avoids with its inherent sliding prevention.
IX Conclusion
@@
We have introduced a novel sim2real robot control strategy capable of robustly folding sheets of paper of varying materials and geometries along symmetrical centerlines with only a single manipulator. Our framework incorporates a combination of techniques spanning several disciplines, including physical simulation, machine learning, scaling analysis, and path planning. The effectiveness of the framework was demonstrated through extensive real world experiments against both natural and state-of-the-art paper folding strategies. Furthermore, an efficient, near real-time visual feedback algorithm was implemented that further minimizes folding error. Our closed-loop model-predictive control algorithm successfully accomplished challenging scenarios such as folding stiff cardboard with highly consistent accuracy.
@@
In future work, we hope to tackle the difficult problem of creating arbitrary creases along sheets of paper with non-symmetric centerlines. Such non-symmetric paper sheets can no longer be represented as a reduced-order 2D elastic rod model, hence requiring a more sophisticated shell-based formulation. Additionally, precisely folding paper with preexisting creases and folds will be a crucial step to accomplishing elaborate folding tasks, such as robotic origami. We believe that our optimal symmetrical folding trajectories can serve as a valuable “seed” or initial guess when optimizing for asymmetrical folds. Moving forward, we anticipate exploring solutions that take advantage of generalized problem formulations with data-driven control schemes, such as reinforcement learning.